A computer has a bus with a 8 nsec cycle time, during which it can read or write a 64-bit word from memory. The computer has a disk that uses the bus and transfers 160 Mbytes/sec. The CPU normally fetches and executes one 64-bit instruction every 1 nsec. How much does the disk slow down the CPU?
Answer #1

64 bit can read by computer in 8

160 mB can read in 8/64*10^6sec=125000*160=2*10^7

the cpu runs at 1n sec

time=1/1nsec

time=10^9;

10^9-2*10^7=0.98*10^9sec(slow down)

so 0.98/1*100=98% it means it slow down by 2%